# A Survey of Vibe Coding with Large Language Models

### v1

@@ 1 Introduction
Large language models (LLMs) have significantly advanced artificial intelligence through conversational systems capable of fluent natural language understanding and generation [1, 2]. Early adoption in software development positioned LLMs as supplementary assistants—developers employed natural language prompts to generate code snippets, but significant accuracy limitations necessitated manual review and iterative debugging throughout the software development lifecycle [3, 4, 5, 6, 7].
@@
The emergence of advanced architectures like GPT-4 [8] and Claude Sonnet 4 [9] enabled qualitative improvements, leading to Coding Agents capable of autonomously completing programming tasks through dynamic environmental interaction via shell commands, file operations, and test execution [10]. These agents have demonstrated rapid progress on real-world programming tasks. Taking SWE-bench as an example [11], SWE-agent reached 12.5% with custom interfaces [12], AutoCodeRover achieved 19.0% resolution through code search and fault localization [13], Agentless attained 27.3% [14], OpenHands achieved 53% on SWE-bench Verified [15], and self-improving agents demonstrated 17–53% performance gains on SWE-bench Verified [16].
@@
With the advancement of large language models such as GPT-5 Pro [17] and Claude Sonnet 4.5 [9], llm-based coding agent capabilities have achieved significant breakthroughs, giving rise to “Vibe Coding”—a paradigm where developers rely on AI-generated code without line-by-line inspection, engaging instead in iterative cycles of natural language requirement articulation, execution observation, and feedback [18, 19, 20, 21]. Coding Agents go beyond code generation—they autonomously configure environments, execute programs, self-diagnose errors, and update implementations. This represents a substantial elevation in human trust and a departure from traditional comprehension mandates toward outcome-oriented validation [22, 23, 24, 25].
@@
However, possessing powerful agents proves insufficient. Task complexity exposes fundamental limitations in unstructured natural language instructions, which fail to convey nuanced requirements and architectural constraints [11, 26]. Empirical evidence reveals that experienced developers using Cursor with Claude experienced 19% increased completion time rather than anticipated productivity gains [27]. Effective human-AI collaboration demands systematic prompt engineering and context engineering [28], structured instructions [29, 3, 30, 31], and balanced agency distribution across various distinct interaction types [32, 33].
@@
To address this critical gap, this survey provides the first comprehensive and systematic review of Vibe Coding with Large Language Models. As shown in Figure 2, we introduce Vibe Coding as a dynamic triadic relationship among human developers, software projects, and coding agents, providing its first formal definition as an engineering discipline through a Constrained Markov Decision Process [34]. Building upon this theoretical foundation, we distill Vibe Coding workflows into five development models—Unconstrained Automation Model, Iterative Conversational Collaboration Model, Planning-Driven Model, Test-Driven Model, and Context-Enhanced Model—representing the first comprehensive synthesis of existing practices. Through this framework, we: (1) establish rigorous theoretical foundations for understanding human-agent collaboration in software development; (2) provide developers with actionable guidance for selecting and implementing appropriate development strategies; and (3) identify critical challenges and future directions spanning technical infrastructure, security mechanisms, and human factors. This work serves as both a conceptual foundation for the emerging field of AI-augmented software engineering and a technical roadmap for advancing research and practice in coding agent systems.
@@
Vibe Coding Large Language Models for Coding (§ 4) Data Foundation of Code LLMs (§ 4.1) Codex [35], MathPile [36], Arctic-SnowCoder [37], OpenCoder [38], SwallowCode [39], WizardCoder [5], OctoPack [40], OpenCodeInstruct [41], CodeArena [42], CodeUltraFeedback [43], Self-Instruct [44], etc. Pre-training Techniques (§ 4.2) BERT [45], T5 [46], CodeBPE [47], GraphCodeBERT [48], CodeT5 [49], Code Llama [7], Birdie [50], Agent-Q [51], GPT-3 [1], CodeBERT [52], UniXcoder [53], CoDist [54], NatGen [55], PLBART [56], CodeT5+ [57], etc. Post-training Techniques (§ 4.3) VeriCoder [58], Flan-PaLM [59], FLAN [60], SparkRA [61], PanGu-Coder2 [62], LoRA [63], Adapter Modules [64], LIMA [65], AlpaGasus [66], FinDPO [67], DPO [68], RAG-Gym [69], RLHF [70], DUMP [71], etc. LLM-based Coding Agent (§ 5) Planning and Decomposition Capability (§ 5.1) Chain-of-Thought [72], AgentGen [73], ReAct [74], DPPM [75], HuggingGPT [76], Zero-shot-CoT [77], Auto-CoT [78], Tree of Thoughts [79], HyperTree Planning [80], CodePlan [81], SQLucid [82], HQD [83], etc. Memory Mechanism (§ 5.2) Transformer [84], MMS [85], A-MEM [86], MemEngine [87], TaSL [88], MemoryBank [89], Generative Agents [90], Zero-Shot Planner [91], RET-LLM [92], etc. Action Execution (§ 5.3) Toolformer [93], Large Knowledge Model [94], Alpha-UMi [95], MCP [96], MCP-Zero [97], AutoTools [98], OpenHands [15], CodeAgent [99], RAIT [100], ScaleMCP [101], Doc2Agent [102], SE-Agent [103], AgentCoder [104], etc. Reflection (§ 5.4) Self-Refine [105], Saarthi [106], Repeton [107], Self-Planning [108], PairCoder [109], Self-Debugging [110], Chained [111], TiCoder [112], ITDCG [113], Self-critique [114], N2M-RSI [115], ReVeal [116], ProCoder [117], RGD [118], etc. Agent Collaboration (§ 5.5) CoMAL [119], DRF [120], TeamMedAgents [121], AgentMesh [122], Hybrid [123], LMA [124], MASs [125], Agent Forest [126], Chatdev [10], MapCoder [127], etc. Development Environment (§ 6) Isolated Execution Runtime (§ 6.1) Virtual Earth Cloud [128], KUNERVA [129], FunDa [130], COCOS [131], SCHEMA lab [132], Singularity [133], DRIVE [134], MultiPL-E [135], TableGPT2 [136], SWE-bench [11], SandboxEval [137], AutoSafeCoder [138], Secure SDLC [139], etc. Interactive Development Interface (§ 6.2) Designing PairBuddy [140], MultiMind [141], Language server protocol [142], Git Context Controller [143], StackSpot AI [144], LSP [145], Opik [146], etc. Distributed Orchestration Platform (§ 6.3) TosKer [147], TOSCAdata [148], TORCH [149], AutoGen [150], Autogen [151], AUTOGEN [152], CrewAI [153], MetaGPT [154], Microservices [155], XP [156], FogArm [157], Agentic [158], LLMOps [159], PharmaSwarm [160], EGI [161], etc. Feedback Mechanisms (§ 7) Compiler Feedback (§ 7.1) RLCF [162], CYCLE [163], AlphaTrans [164], RLEF [165], AceCoder [166], VisCoder [167], InterCode [26], The Art of Repair [168], Rtlfixer [169], etc. Execution Feedback (§ 7.2) PerfCodeGen [170], ExecutionAgent [171], TypeTest [172], ProjectTest [173], PyCapsule [174], ASTER [175], CodeT [176], STROT [177], KPC [178], APT [179], Seeker [180], TestGen-LLM [181], Healer [182], Agentic workflow [183], etc. Human Feedback (§ 7.3) Grounded Copilot [184], Cutting-Search [185], Deep Sets [186], MA-RLHF [187], ClarifyGPT [188], PPO [189], RLSF [190], CAGSR–vLLM–MTC [191], etc. Self-Refinement Feedback (§ 7.4) CRITIC [192], BioAgents [193], Self-Organized Agents [194], Multi-Agent Collaboration [195], LLM-Agent-UMF [196], SWE-Search [197], N-Critics [198], Self-Review Framework [199], SAMULE [200], Reflexion [201], etc.
@@ 2.1 Related Surveys
Several surveys have documented the evolution of large language models, covering LLM architectures, training paradigms, and capabilities [202, 203]. These works examine Transformer variants with a focus on efficient architectures and long-context capabilities [204, 205, 206]. The historical trajectory from BERT to ChatGPT traces the development of foundation models, examining both opportunities and risks [207]. Specialized surveys address evaluation methodologies, efficiency from model-centric and data-centric perspectives, and specific capabilities including text generation and knowledge-enhanced PLMs [208].
@@
Building on these architectural foundations, research has turned to techniques for utilizing pre-trained models without additional training. Prompt engineering and in-context learning have emerged as fundamental techniques, with taxonomies covering extensive prompting methods, applications, security considerations, and performance across Natural Language Processing (NLP) tasks [209, 210, 211]. In-context learning mechanisms are explored with context engineering emerging as a formal discipline [212, 28]. Chain-of-Thought (CoT) reasoning has proven particularly effective, with taxonomies examining Chain-of-X paradigms and investigating long CoT and multimodal CoT reasoning [213, 214, 215]. Multimodal large language models represent a rapidly advancing frontier, with surveys examining architectures, training methods, and vision-language integration across multiple data modalities [216, 217, 218, 219].
@@
When in-context learning proves insufficient, post-training methods offer pathways to align models with specific requirements and enhance reasoning capabilities. Reinforcement learning approaches, including Proximal Policy Optimization (PPO), Q-Learning, and Actor-Critic methods, have been surveyed with particular emphasis on Reinforcement Learning from Human Feedback (RLHF), Reinforcement Learning from AI Feedback (RLAIF), and Direct Preference Optimization (DPO) [220, 221, 222, 223, 224]. Instruction tuning and supervised fine-tuning methodologies are reviewed covering dataset construction and training strategies, with examinations of data selection approaches for enhanced instruction-following capabilities [225]. Alignment research categorizes methods into outer and inner alignment with adversarial considerations, while exploring training-free alignment and personalized alignment techniques [226]. DPO emerges as an reinforcement learning RL-free alternative to RLHF, with taxonomies categorizing data strategies, learning frameworks, and constraint mechanisms [227]. Post-training paradigms are explored covering fine-tuning, alignment, reasoning, efficiency, and domain adaptation, with parameter-efficient methods including Low-Rank Adaptation (LoRA) and adapters providing experimental comparisons of computational overhead [228].
@@
The integration of tool use and planning capabilities transforms LLMs from passive models into active agents. Foundational surveys establish frameworks covering agent construction, brain-perception-action architectures, and autonomous decision-making capabilities, while providing unified taxonomies across benchmarks spanning reasoning and code generation [229, 230, 231]. Multi-agent systems are examined covering agent profiling, communication protocols, and collaborative workflows across complex task-solving scenarios [232, 233]. Agent capabilities are addressed through specialized surveys: tool use with Retrieval-Augmented Generation (RAG) and feedback learning [234, 235], planning mechanisms including task decomposition and memory [221], single and multi-agent architectures for reasoning and tool execution, and memory mechanisms with short-term and long-term analysis [236]. Evaluation methodologies cover planning, tool use, self-reflection, and application-specific benchmarks [237, 238]. Domain-specific applications span web automation [239], scientific discovery across life sciences and materials science [240], Operating System agents with GUI interaction [241], and self-evolving agents with feedback loops and lifelong learning capabilities [242]. Of particular relevance to this work, recent surveys examine coding agents with single and multi-agent architectures across the software development lifecycle, covering planning, context management, and tool integration with benchmarking frameworks [243, 244].

### v2

@@ 1 Introduction
Large language models (LLMs) have significantly advanced artificial intelligence through conversational systems capable of fluent natural language understanding and generation [brown2020language, ouyang2022training]. Early adoption in software development positioned LLMs as supplementary assistants—developers employed natural language prompts to generate code snippets, but significant accuracy limitations necessitated manual review and iterative debugging throughout the software development lifecycle [zhang2023repocoder, nijkamp2023codegen, luo2023wizardcoder, li2023starcoder, roziere2023code].
@@
The emergence of advanced architectures like GPT-4 [achiam2023gpt] and Claude Sonnet 4 [anthropic2025claude45] enabled qualitative improvements, leading to Coding Agents capable of autonomously completing programming tasks through dynamic environmental interaction via shell commands, file operations, and test execution [qian2023chatdev]. These agents have demonstrated rapid progress on real-world programming tasks. Taking SWE-bench as an example [jimenez2023swe], SWE-agent reached 12.5% with custom interfaces [yang2024swe], AutoCodeRover achieved 19.0% resolution through code search and fault localization [zhang2024autocoderover], Agentless attained 27.3% [xia2024agentless], OpenHands achieved 53% on SWE-bench Verified [wang2024openhands], and self-improving agents demonstrated 17–53% performance gains on SWE-bench Verified [robeyns2025self].
@@
With the advancement of large language models such as GPT-5 Pro [openai2025gpt5] and Claude Sonnet 4.5 [anthropic2025claude45], llm-based coding agent capabilities have achieved significant breakthroughs, giving rise to “Vibe Coding”—a paradigm where developers rely on AI-generated code without line-by-line inspection, engaging instead in iterative cycles of natural language requirement articulation, execution observation, and feedback [karpathy2025, ray2025review, sapkota2025vibe, horvat2025vibe]. Coding Agents go beyond code generation—they autonomously configure environments, execute programs, self-diagnose errors, and update implementations. This represents a substantial elevation in human trust and a departure from traditional comprehension mandates toward outcome-oriented validation [novikov2025alphaevolve, ho2025verilogcoder, zhao2025mage, zeng2025glm].
@@
However, possessing powerful agents proves insufficient. Task complexity exposes fundamental limitations in unstructured natural language instructions, which fail to convey nuanced requirements and architectural constraints [jimenez2023swe, yang2023intercode]. Empirical evidence reveals that experienced developers using Cursor with Claude experienced 19% increased completion time rather than anticipated productivity gains [becker2025measuring]. Effective human-AI collaboration demands systematic prompt engineering and context engineering [mei2025survey], structured instructions [nashid2023retrieval, zhang2023repocoder, zhou2023large, white2023prompt], and balanced agency distribution across various distinct interaction types [treude2025developers, fragiadakis2024evaluating].
@@
To address this critical gap, this survey provides the first comprehensive and systematic review of Vibe Coding with Large Language Models. As shown in Figure 2, we introduce Vibe Coding as a dynamic triadic relationship among human developers, software projects, and coding agents, providing its first formal definition as an engineering discipline through a Constrained Markov Decision Process [altman2021constrained]. Building upon this theoretical foundation, we distill Vibe Coding workflows into five development models—Unconstrained Automation Model, Iterative Conversational Collaboration Model, Planning-Driven Model, Test-Driven Model, and Context-Enhanced Model—representing the first comprehensive synthesis of existing practices. Through this framework, we: (1) establish rigorous theoretical foundations for understanding human-agent collaboration in software development; (2) provide developers with actionable guidance for selecting and implementing appropriate development strategies; and (3) identify critical challenges and future directions spanning technical infrastructure, security mechanisms, and human factors. This work serves as both a conceptual foundation for the emerging field of AI-augmented software engineering and a technical roadmap for advancing research and practice in coding agent systems.
@@
Vibe Coding Large Language Models for Coding (§ 4) Data Foundation of Code LLMs (§ 4.1) Codex [chen2021evaluating], MathPile [wang2023mathpile], Arctic-SnowCoder [wei2024arctic], OpenCoder [huang2024opencoder], SwallowCode [fujii2025rewriting], WizardCoder [luo2023wizardcoder], OctoPack [muennighoff2023octopack], OpenCodeInstruct [ahmad2025opencodeinstruct], CodeArena [yang2024evaluating], CodeUltraFeedback [weyssow2024codeultrafeedback], Self-Instruct [wang2022self], etc. Pre-training Techniques (§ 4.2) BERT [devlin2019bert], T5 [raffel2019exploring], CodeBPE [chirkova2023codebpe], GraphCodeBERT [guo2020graphcodebert], CodeT5 [wang2021identifier], Code Llama [roziere2023code], Birdie [blouir2024birdie], Agent-Q [jern2025agent], GPT-3 [brown2020language], CodeBERT [feng2020codebert], UniXcoder [guo2022unixcoder], CoDist [huang2023program], NatGen [chakraborty2022natgen], PLBART [ahmad2021unified], CodeT5+ [wang2023open], etc. Post-training Techniques (§ 4.3) VeriCoder [wei2025vericoder], Flan-PaLM [chung2022scaling], FLAN [wei2021finetuned], SparkRA [wu2024sparkra], PanGu-Coder2 [shen2023pangu], LoRA [hu2022lora], Adapter Modules [houlsby2019parameter], LIMA [zhou2023lima], AlpaGasus [chen2023alpagasus], FinDPO [iacovides2025findpo], DPO [rafailov2023direct], RAG-Gym [xiong2025rag], RLHF [stiennon2020learning], DUMP [wang2025dump], etc. LLM-based Coding Agent (§ 5) Planning and Decomposition Capability (§ 5.1) Chain-of-Thought [wei2022chain], AgentGen [hu2024agentgen], ReAct [yao2022react], DPPM [lu2025decompose], HuggingGPT [shen2023hugginggpt], Zero-shot-CoT [kojima2022large], Auto-CoT [zhang2022automatic], Tree of Thoughts [yao2023tree], HyperTree Planning [gui2025hypertree], CodePlan [wen2024codeplan], SQLucid [tian2024sqlucid], HQD [patel2022question], etc. Memory Mechanism (§ 5.2) Transformer [vaswani2017attention], MMS [zhang2025multiple], A-MEM [xu2025mem], MemEngine [zhang2025memengine], TaSL [feng2024tasl], MemoryBank [zhong2023memorybank], Generative Agents [park2023generative], Zero-Shot Planner [huang2022language], RET-LLM [modarressi2023ret], etc. Action Execution (§ 5.3) Toolformer [schick2023toolformer], Large Knowledge Model [chen2023large], Alpha-UMi [shen2024small], MCP [hou2025model], MCP-Zero [fei2025mcp], AutoTools [shi2024tool], OpenHands [wang2024openhands], CodeAgent [zhang2024codeagent], RAIT [srinivas2024retrieval], ScaleMCP [lumer2025scalemcp], Doc2Agent [ni2025doc2agent], SE-Agent [lin2025agent], AgentCoder [huang2023agentcoder], etc. Reflection (§ 5.4) Self-Refine [madaan2023self], Saarthi [kumar2025saarthi], Repeton [vinh2025repeton], Self-Planning [jiang2023self], PairCoder [zhang2024pair], Self-Debugging [chen2023teaching], Chained [ashrafi2025enhancing], TiCoder [fakhoury2024llm], ITDCG [lahiri2022interactive], Self-critique [ho2025self], N2M-RSI [ando2025when], ReVeal [jin2025reveal], ProCoder [bi2024iterative], RGD [jin2024rgd], etc. Agent Collaboration (§ 5.5) CoMAL [yao2024comal], DRF [lou2025drf], TeamMedAgents [mishra2025teammedagents], AgentMesh [khanzadeh2025agentmesh], Hybrid [williams2025multi], LMA [he2025llm], MASs [tran2025multi], Agent Forest [li2024more], Chatdev [qian2023chatdev], MapCoder [islam2024mapcoder], etc. Development Environment (§ 6) Isolated Execution Runtime (§ 6.1) Virtual Earth Cloud [santoro2023virtual], KUNERVA [lee2023kunerva], FunDa [lounissi2025funda], COCOS [baresi2020cocos], SCHEMA lab [adamidi2025virtual], Singularity [kurtzer2017singularity], DRIVE [zhou2022drive], MultiPL-E [cassano2023multipl], TableGPT2 [su2024large], SWE-bench [jimenez2023swe], SandboxEval [rabin2025sandboxeval], AutoSafeCoder [nunez2024autosafecoder], Secure SDLC [gajbhiye2024secure], etc. Interactive Development Interface (§ 6.2) Designing PairBuddy [robe2022designing], MultiMind [donato2025multimind], Language server protocol [bork2023language], Git Context Controller [wu2025git], StackSpot AI [pinto2023developer], LSP [gunasinghe2021language], Opik [koc2025mind], etc. Distributed Orchestration Platform (§ 6.3) TosKer [brogi2018tosker], TOSCAdata [dehury2021toscadata], TORCH [tomarchio2021torch], AutoGen [wu2023autogen], Autogen [zhu2023autogen], AUTOGEN [porsdam2023autogen], CrewAI [taulli2025crewai], MetaGPT [zhou2024metagpt], Microservices [singh2023microservices], XP [beck2000extreme], FogArm [bisicchia2023continuous], Agentic [joshi2025architecting], LLMOps [choi2025intelligent], PharmaSwarm [song2025llm], EGI [zeng2025edge], etc. Feedback Mechanisms (§ 7) Compiler Feedback (§ 7.1) RLCF [jain2023coarse], CYCLE [ding2024cycle], AlphaTrans [ibrahimzada2024alphatrans], RLEF [gehring2024rlef], AceCoder [zeng2025acecoder], VisCoder [ni2025viscoder], InterCode [yang2023intercode], The Art of Repair [ruiz2025art], Rtlfixer [tsai2023rtlfixer], etc. Execution Feedback (§ 7.2) PerfCodeGen [peng2024perfcodegen], ExecutionAgent [bouzenia2024you], TypeTest [liu2025llm], ProjectTest [wang2025projecttest], PyCapsule [adnan2025large], ASTER [pan2025aster], CodeT [chen2022codet], STROT [rath2025structured], KPC [ren2023from], APT [zhang2024llm], Seeker [zhang2024seeker], TestGen-LLM [alshahwan2024automated], Healer [sun2024llm], Agentic workflow [schultz2025potential], etc. Human Feedback (§ 7.3) Grounded Copilot [barke2023grounded], Cutting-Search [marozzo2025iterative], Deep Sets [zaheer2017deep], MA-RLHF [chai2024ma], ClarifyGPT [mu2024clarifygpt], PPO [schulman2017proximal], RLSF [niekerk2025post], CAGSR–vLLM–MTC [kiruluta2025history], etc. Self-Refinement Feedback (§ 7.4) CRITIC [gou2023critic], BioAgents [mehandru2025bioagents], Self-Organized Agents [ishibashi2024self], Multi-Agent Collaboration [talebirad2023multi], LLM-Agent-UMF [hassouna2024llm], SWE-Search [antoniades2024swe], N-Critics [mousavi2023critics], Self-Review Framework [park2025self], SAMULE [ge2025samule], Reflexion [shinn2023reflexion], etc.
@@ 2.1 Related Surveys
Several surveys have documented the evolution of large language models, covering LLM architectures, training paradigms, and capabilities [minaee2024large, zheng2023survey]. These works examine Transformer variants with a focus on efficient architectures and long-context capabilities [tay2022efficient, wan2023efficient, liu2025comprehensive]. The historical trajectory from BERT to ChatGPT traces the development of foundation models, examining both opportunities and risks [zhou2024comprehensive]. Specialized surveys address evaluation methodologies, efficiency from model-centric and data-centric perspectives, and specific capabilities including text generation and knowledge-enhanced PLMs [chang2024survey].
@@
Building on these architectural foundations, research has turned to techniques for utilizing pre-trained models without additional training. Prompt engineering and in-context learning have emerged as fundamental techniques, with taxonomies covering extensive prompting methods, applications, security considerations, and performance across Natural Language Processing (NLP) tasks [schulhoff2024prompt, sahoo2024systematic, vatsal2024survey]. In-context learning mechanisms are explored with context engineering emerging as a formal discipline [dong2022survey, mei2025survey]. Chain-of-Thought (CoT) reasoning has proven particularly effective, with taxonomies examining Chain-of-X paradigms and investigating long CoT and multimodal CoT reasoning [chu2023navigate, xia2024beyond, wang2025multimodal]. Multimodal large language models represent a rapidly advancing frontier, with surveys examining architectures, training methods, and vision-language integration across multiple data modalities [yin2024survey, zhang2024mm, liang2024comprehensive, wu2023multimodal].
@@
When in-context learning proves insufficient, post-training methods offer pathways to align models with specific requirements and enhance reasoning capabilities. Reinforcement learning approaches, including Proximal Policy Optimization (PPO), Q-Learning, and Actor-Critic methods, have been surveyed with particular emphasis on Reinforcement Learning from Human Feedback (RLHF), Reinforcement Learning from AI Feedback (RLAIF), and Direct Preference Optimization (DPO) [srivastava2025technical, zhang2025survey, cao2024survey, kaufmann2024survey, ma2023training]. Instruction tuning and supervised fine-tuning methodologies are reviewed covering dataset construction and training strategies, with examinations of data selection approaches for enhanced instruction-following capabilities [shengyu2023instruction]. Alignment research categorizes methods into outer and inner alignment with adversarial considerations, while exploring training-free alignment and personalized alignment techniques [tie2025survey]. DPO emerges as an reinforcement learning RL-free alternative to RLHF, with taxonomies categorizing data strategies, learning frameworks, and constraint mechanisms [xiao2024comprehensive]. Post-training paradigms are explored covering fine-tuning, alignment, reasoning, efficiency, and domain adaptation, with parameter-efficient methods including Low-Rank Adaptation (LoRA) and adapters providing experimental comparisons of computational overhead [xu2023parameter].
@@
The integration of tool use and planning capabilities transforms LLMs from passive models into active agents. Foundational surveys establish frameworks covering agent construction, brain-perception-action architectures, and autonomous decision-making capabilities, while providing unified taxonomies across benchmarks spanning reasoning and code generation [su2024language, zhao2023depth, ferrag2025llm]. Multi-agent systems are examined covering agent profiling, communication protocols, and collaborative workflows across complex task-solving scenarios [guo2024large, chen2024survey]. Agent capabilities are addressed through specialized surveys: tool use with Retrieval-Augmented Generation (RAG) and feedback learning [huang2024understanding, masterman2024landscape], planning mechanisms including task decomposition and memory [zhang2025survey], single and multi-agent architectures for reasoning and tool execution, and memory mechanisms with short-term and long-term analysis [xu2025llm]. Evaluation methodologies cover planning, tool use, self-reflection, and application-specific benchmarks [yehudai2025survey, mohammadi2025evaluation]. Domain-specific applications span web automation [ning2025survey], scientific discovery across life sciences and materials science [wei2025ai], Operating System agents with GUI interaction [hu2025agents], and self-evolving agents with feedback loops and lifelong learning capabilities [fang2025comprehensive]. Of particular relevance to this work, recent surveys examine coding agents with single and multi-agent architectures across the software development lifecycle, covering planning, context management, and tool integration with benchmarking frameworks [dong2025survey, wang2025ai].